myininaya (myininaya):

@utkFresh \[\int \sin(ax) dx \] Let u=ax So du=a dx du/a=dx So we have \[\int \sin(u) \frac{du}{a} \] \[\frac{1}{a} \int \sin(u) du \\ \frac{1}{a} (-cos(u))+C \\ frac{-1}{a}cos(ax)+C

OpenStudy (anonymous):

\[\frac{1}{a} \int \sin(u) du \\ \frac{1}{a} (-\cos(u))+C \\ \frac{-1}{a}\cos(ax)+C \]

myininaya (myininaya):

\[\frac{1}{a} \int \sin(u) du \\ \frac{1}{a} (-cos(u))+C \\\ \frac{-1}{a}\cos(ax)+C \]

OpenStudy (anonymous):

no charge

myininaya (myininaya):

no I fixed it !

myininaya (myininaya):

I want a refund.

TheSmartOne (thesmartone):

??? I am so confused.

TheSmartOne (thesmartone):

Maybe its because I don't know Calculus...

myininaya (myininaya):

I just realized someone messaged me 26 minutes ago and they wanted to know if there is a formula. a does not equal 0

TheSmartOne (thesmartone):

oh yeah because then it would be undefined...